OpenStudy (anonymous):

find the slope of the line tangent to the following curve at x=1...

OpenStudy (anonymous):

\[y = \sin \left[ x - \tan(\frac{ \Pi }{ 4 }x ^{66}) \right] + x ^{\frac{ 1 }{ 1+66\Pi }}\]

OpenStudy (anonymous):

slope of the line TANGENT to this curve at x =1

OpenStudy (anonymous):

answer: -99Pi

OpenStudy (anonymous):

\[y' = \cos \left[ x - \tan(\frac{ \Pi }{ 4 }x ^{66})\right](1 - \sec ^{2}(\frac{ \Pi }{ 4 }x ^{66}))(2\sec(\frac{ \Pi }{ 4 }x ^{66}))(\frac{ 33\Pi }{ 2 }x ^{65}) + \frac{ 1 }{1+66\Pi }x ^{\frac{ 1 }{ 1+66\Pi }-1}\]

OpenStudy (anonymous):

is what i get for the derivative

OpenStudy (anonymous):

damn even this is too long

OpenStudy (anonymous):

could you guys verify the derivative expression

OpenStudy (anonymous):

if youre working on this

OpenStudy (anonymous):

what kind of sadistic math teacher gave you this problem....

OpenStudy (helder_edwin):

what is the power of x to the right??

OpenStudy (anonymous):

you have to use chain rule then substitute x=1.

OpenStudy (anonymous):

your y(prime) is correct up to 2sec...

OpenStudy (anonymous):

\[66pix^{65}/4\]

OpenStudy (anonymous):

the above should be in place of your 2sec value... then add the derivative of that last value and substitute x=1